# E-scooter effects on public transport demand: a case study in Santiago, Chile

### v1

@@ 4 Methodology
To examine the impact of scooters, we conducted a study utilizing the difference-in-differences (Diff-in-Diff) approach. Our spatial unit of analysis consisted of the geographical zones defined by the DTPM. The urban areas (geographical zones) were divided into three distinct groups using a clustering method to account for spatial heterogeneity and ensure that comparisons between control and treatment zones within each cluster were valid. In this section, we present the details of the variable selection process for clustering and the application of the Diff-in-Diff methodology.
@@
We compiled an initial list of variables that are typically considered to influence travel behavior and assessed their relevance in the context of Santiago and this study. The selected variables include sociodemographic variables from the 2017 census, such as underage population, working population, retired population, percentage of female population, average educational level, and population density. Additionally, we included variables from ADATRAP dataset to capture travel behavior, such as the average number of public transport trips per person, average time of the first public transport trip made each day, and average public transport trip distance. We also included variables to determine each zone’s accessibility to public transport, mapping the metro network within Santiago and counting the number of metro stations and bus stops per zone using data from OpenStreetMap and the General Transit Feed Specification (GTFS). These indicators were calculated for each zone in our design. A comprehensive list of these variables can be found in Table 2. Additionally, Figure 5 and Figure 6 show the distribution of these variables across the different zones of the city. We can see that there’s a central point in the city that is highly densely populated, and as we get farther away from it the population spreads out.
@@
Table 2: Summary of Zone Clustering Variables. This table presents a comprehensive summary of the variables employed in our study. Each column includes the variable name, a brief description, the data source (represented by a number), the role of the variable in the analysis (use), its unit of measurement, and key statistical indicators such as the mean and standard deviation. The variables span demographic information (e.g., underage population, working population), educational level, and transport-related metrics (e.g., number of bus stops, bus stop density). The data sources are as follows: (1) 2017 Census, (2) OpenStreetMap, and (3) ADATRAP.
@@
We started by using Moran’s I to analyze the spatial autocorrelation of variables from the previous section. This method helped identify variables with significant spatial patterns, ensuring that the selected variables were not randomly distributed across space but had meaningful spatial dependencies. A positive Moran’s I value indicates clustering over space, while a negative Moran’s I indicates dispersion. We chose 0.25 as a threshold for qualifying variables. Additionally, we established a selection criterion based on the p-sim value of each variable, representing the likelihood of the observed values being randomly spatially distributed. We set 0.05 as the upper limit for the p-sim value of each variable.
@@
After applying the clustering algorithm and dividing the zones into three groups of regions, we excluded 56 zones from the study despite them having no origin and destination trips involving scooters, as they could still experience mobility impacts post-intervention. The criteria for exclusion were zones intersecting with a 1,440 m buffer applied to each scooter trajectory, a distance chosen based on the average trip distance calculated from our data. Figure 11 displays the control zones, treatment zones, and the excluded zones. Figure 12, along with Table 3, presents the distribution and detailed breakdown of control and treatment zones by cluster, excluding the 56 mentioned zones.
@@
Finally, we measured how the implementation of e-scooters affects public transportation demand. Given the number of public transportation trips often exhibits significant variability and over-dispersion, meaning that the variance exceeds the mean. This makes traditional Poisson regression unsuitable for our dataset. Instead, we applied two distinct Negative Binomial NB regression models (Cameron and Trivedi, 1986): one for trips generation (trips originating from public transport stations) and one for trips attraction (trips ending at public transport stations). The NB models are well-suited to handle over-dispersed count data and allow us to better capture the variability in public transportation trips across different zones, providing a more accurate understanding of the factors influencing both trip generation and attraction.
The general form of the NB model is defined as follows:
$f(y_{i}\mid\mathbf{x}_{i})=\frac{\Gamma(y_{i}+\alpha)}{y_{i}!\Gamma(\alpha)}\left(\frac{\alpha}{\alpha+u_{i}}\right)^{\alpha}\left(\frac{u_{i}}{\alpha+u_{i}}\right)^{y_{i}},$
where $y_{i}$ is the response variable, i.e., the count of public transportation trips in Traffic Analysis Zones (TAZ), $\mathbf{x}_{i}$ is the set of independent variables, $\Gamma(\cdot)$ is the gamma function, and $\alpha$ is the dispersion parameter. The conditional mean and variance are given by:
$\mathbf{E}(y_{i}\mid\mathbf{x}_{i})=u_{i}=\exp(\mathbf{x}_{i}^{T}\mathbf{\beta});\text{Var}(y_{i}\mid\mathbf{x}_{i})=u_{i}\left(1+\frac{1}{\alpha}u_{i}\right).$
@@
This choice of model is coherent with studies that have applied similar approaches to transportation demand across various modes  (e.g., Schimohr and Scheiner, 2021; Hu et al., 2022; Manville et al., 2023), and with previous research analyzing behavioral and spatial patterns in Santiago  (Graells-Garrido et al., 2017; Beiró et al., 2018; Graells-Garrido et al., 2023b).
Here, the set of independent variables that define each trip count contains the following:
  • 1.
    year: a binary variable which has two possible values: 2019 and 2018, with 2018 as reference value;
  • 2.
    type of day: a categorical variable with four possible values: monday-to-thursday, friday, saturday, and sunday. monday-to-thursday is used as the reference category.
  • 3.
    time period: a categorical variable that describes the time of the day when the trip was performed, including the following values: morning valley, morning peak, lunch, afternoon valley, afternoon peak, night, and night valley, with morning valley as reference value. The time periods are defined in the table provided in the Appendix A (see Table A.6);
  • 4.
    scooter trips origins: a dummy variable where 1 indicates that there are e-scooter trips originating in the specific TAZ, and 0 indicates none;
  • 5.
    scooter trips destinations: a dummy variable where 1 indicates that there are e-scooter trips ending in the specific TAZ, and 0 indicates none;
  • 6.
    working population: refers to the log-transformed count of the working population that resides in the specific TAZ;
  • 7.
    relative educational level: denotes the average educational level for the specific TAZ measured as years of education, relative to the regional average;
  • 8.
    metro: denotes the number of metro lines within a distance of less than 750 meters from the specific TAZ area.
Using these independent variables, we define the negative binomial regression models for both generation and attraction of trips. For the generation model, which examines trips originating from public transport stations, we use the following expression:
$\begin{split}\ln\left({\textbf{E}}[y_{\text{gen}}\mid\mathbf{x}_{\text{gen}}]\right)=&\beta_{0}+\beta_{1}\text{year}+\beta_{2}\text{type\_of\_day}+\beta_{3}\text{time\_period }+\\ &\beta_{4}\text{scooter\_trips\_destinations}+\beta_{5}\text{working\_population }+\\ &\beta_{6}\text{relative\_educational\_level}+\beta_{7}\text{metro}\end{split}$
Similarly, for the attraction model, which examines trips ending at public transport stations, the model is expressed as:
$\begin{split}\ln\left({\textbf{E}}[y_{\text{att}}\mid\mathbf{x}_{\text{att}}]\right)=&\beta_{0}+\beta_{1}\text{year}+\beta_{2}\text{type\_of\_day}+\beta_{3}\text{time\_period }+\\ &\beta_{4}\text{scooter\_trips\_origins}+\beta_{5}\text{working\_population }+\\ &\beta_{6}\text{relative\_educational\_level}+\beta_{7}\text{metro}\end{split}$ (1)
The key difference between these models is in the scooter trip variable: the generation model $f(y_{gen}\mid\mathbf{x}_{gen})$ uses scooter trip destinations, while the attraction model $f(y_{att}\mid\mathbf{x}_{att})$ uses scooter trip origins. This distinction allows us to capture the different effects of e-scooters on public transport usage in each scenario.
@@
For each scenario (generation and attraction), we developed three separate models: 1) A model for bus trips only; 2) a model for metro trips only; and 3) a combined model for both bus and metro trips. In all models, the dependent variable is the count of public transport trips in each Traffic Analysis Zone (TAZ). We applied the regression to three regions (Central, Intermediate and Peripheral), each containing its respective control and treatment areas. Our analysis focused on two distinct time periods in fall: the second week of May 2018, and the second week of May 2019. A differences-in-differences approach was utilized to control for temporal trends and other unobservable differences between the zones that implemented scooters (treatment group) and those that did not (control group).
@@ 5 Results
In this section, we present the findings of our study on the impact of shared e-scooters (e-scooters) on public transportation usage in Santiago, Chile. Our analysis, using a negative binomial regression model and a difference-in-differences approach, examines how e-scooter trips affect public transport demand across different urban regions, considering both generation and attraction scenarios. For the metro trip generation and attraction models, we focused on the Central and Intermediate areas, as there are very few metro stations in the Peripheral Region.
@@
Tables in Appendix C show the results of negative binomial models for public transport trip generation, and Table 4 shows the marginal effects of scooter destination zones, located within different regions, on public transport trip generation. In the Central Region, shared e-scooter trips are linked to fewer combined bus and metro trips, with an average reduction of 136.66 trips per hour per zone. This represents a 21.38% reduction, considering the region’s average of 639.26 combined bus and metro trips per hour. However, when focusing specifically on bus trips, shared e-scooters contribute to an increase of 145.22 additional bus trips per hour per zone, reflecting a 76.39% rise from the average of 190.09 bus trips per hour. For metro trips, the impact of e-scooters is not statistically significant.
In the Intermediate Region, which also comprises multiple zones, shared e-scooter trips contribute to an average of 99.77 additional bus and metro trips per hour per zone, leading to a 56.83% increase in combined public transport usage, considering the region’s average of 175.56 trips per hour. The impact on bus trips is more modest, with an increase of 20.10 additional bus trips per hour per zone, corresponding to a 22.52% rise from the region’s average of 89.26 bus trips per hour. For metro trips, e-scooters result in 365.30 additional trips per hour per zone, representing a 70.05% increase relative to the region’s average of 521.46 metro trips per hour.
In the Peripheral Region, shared e-scooter trips generate an average of 45.40 additional bus and metro trips per hour per zone, resulting in a 40.17% increase in combined public transport usage, considering the region’s average of 113.03 trips per hour. The effect on bus trips is more pronounced, with 72.82 additional bus trips per hour per zone, representing an 84.64% increase from the region’s average of 86.04 bus trips per hour.
Mode of Transport Category Central Intermediate Peripheral
Bus or Metro E-scooter impact on trips/hour -136.66 99.77 45.40
Average trips 639.26 175.56 113.03
% Change in trips -21.38% 56.83% 40.17%
Bus E-scooter impact on trips/hour 145.22 20.10 72.82
Average trips 190.09 89.26 86.04
% Change in trips 76.39% 22.52% 84.64%
Metro E-scooter impact on trips/hour - 365.30 -
Average trips - 521.46 -
% Change in trips - 70.05% -
Table 4: Impact of e-scooters on generation of public transport trips. Columns: (1) Mode of transport; (2) Impact of e-scooters on trips; (3) Average number of trips; (4) Percentage change in trips. Data is presented for Central, Intermediate, and Peripheral regions.
@@
Tables in Appendix D show the results of negative binomial models for public transport trip attraction, while Table 5 shows the marginal effects of scooter origin zones, located within different regions, on public transport trip attraction. In the Central Region, e-scooter trips are linked to a large drop in public transport trips ending in these zones. Specifically, in these central zones, e-scooter trips lead to a reduction of approximately 389.17 trips per hour, which corresponds to a 72.06% decline in bus and metro trips, given the average of 540.08 hourly trips in these areas. However, when focusing on bus trips alone, e-scooter trips attract 154.21 additional trips per hour, representing a 38.12% increase over the average of 1186.10 hourly metro trips in central zones. For metro trips, the impact of e-scooters is not statistically significant.
In zones located in the Intermediate Region, e-scooter trips positively influence the attraction of public transport trips. Each e-scooter trip generates an additional 87.17 bus or metro trips per hour, leading to a 63.85% increase over the average of 136.52 hourly trips in these intermediate zones. The impact on bus trips is more modest, with an increase of 14 additional bus trips per hour per zone, corresponding to a 22.29% rise from the region’s average of 62.78 bus trips per hour. Focusing on metro trips specifically, e-scooter trips attract approximately 286.41 trips per hour, resulting in a 62.17% increase compared to the average of 460.68 hourly metro trips in intermediate zones.
In zones located in the Peripheral Region, the impact of e-scooter trips on the attraction of public transport trips is more nuanced. Scooters only have a significant impact on bus trips, attracting 38.62 additional bus trips per hour, representing a notable 77.29% increase compared to the average of 49.97 hourly bus trips in these peripheral zones.
Mode of Transport Category Central Intermediate Peripheral
Bus or Metro E-scooter impact on trips -389.17 87.17
Average trips 540.08 136.52 -
% Change in trips -72.06% 63.85% -
Bus E-scooter impact on trips 51.16 14.00 38.62
Average trips 154.21 62.78 49.97
% Change in trips 38.12% 22.29% 77.29%
Metro E-scooter impact on trips - 286.41 -
Average trips - 460.68 -
% Change in trips - 62.17% -
Table 5: Impact of e-scooters on attraction of public transport trips. Columns: (1) Mode of transport; (2) Impact of e-scooters on trips; (3) Average number of trips; (4) Percentage change in trips. Data is presented for Central, Intermediate, and Peripheral regions.

### v2

@@ 4 Methodology
To examine how e-scooters affect public transport use, we conducted a difference-in-differences analysis using geographical zones defined by DTPM as our spatial units. We divided urban areas into three groups using clustering to account for spatial differences and ensure valid comparisons between control and treatment zones. This section details our variable selection and methodology.
@@
We compiled an initial list of variables that are typically considered to influence travel behavior and assessed their relevance in the context of Santiago and this study. The selected variables include sociodemographic variables from the 2017 census, such as underage population, working population, retired population, percentage of female population, average educational level, and population density. Additionally, we included variables from ADATRAP dataset to capture travel behavior, such as the average number of public transport trips per person, average time of the first public transport trip made each day, and average public transport trip distance. We also included variables to determine each zone’s accessibility to public transport, mapping the metro network within Santiago and counting the number of metro stations and bus stops per zone using data from OpenStreetMap and the General Transit Feed Specification (GTFS). These indicators were calculated for each zone in our design. A comprehensive list of these variables can be found in Table LABEL:table:variables below. Additionally, Figure 5 and Figure 6 show the distribution of these variables across the different zones of the city. We can see that there’s a central point in the city that is highly densely populated, and as we get farther away from it the population spreads out.
Table 1: Summary of Zone Clustering Variables. Each column includes the variable name, a brief description, the data source (represented by a number), the role of the variable in the analysis (use), its unit of measurement, and key statistical indicators such as the mean and standard deviation. The variables span demographic information (e.g., underage population, working population), educational level, and transport-related metrics (e.g., number of bus stops, bus stop density). The data sources are as follows: (1) 2017 Census, (2) OpenStreetMap, and (3) ADATRAP.
@@
We used Moran’s I to analyze spatial autocorrelation and applied two selection criteria: (1) minimum Moran’s I value of 0.25 to ensure meaningful spatial clustering or dispersion, and (2) maximum p-value of 0.05 for statistical significance of the spatial pattern. This ensured that selected variables exhibited statistically significant and substantial spatial dependencies rather than random spatial distribution.
@@
After applying the clustering algorithm and dividing the zones into three groups of regions, we excluded 56 zones from the study. Although these zones had no origin and destination trips involving scooters, they were close to the treatment zones and therefore could still experience indirect mobility impacts post-intervention, potentially contaminating our control group. The criteria for exclusion were zones intersecting with a 1,440 m buffer applied to each scooter trajectory, a distance chosen based on the average trip distance calculated from our data. Figure 11 displays the control zones, treatment zones, and the excluded zones. Figure 12, along with Table 2, presents the distribution and detailed breakdown of control and treatment zones by cluster, excluding the 56 mentioned zones.
@@
Finally, to measure how the implementation of e-scooters affects public transportation demand, we applied a (DiD) approach in conjunction with two distinct Negative Binomial (NB) regression models (Cameron and Trivedi, 1986). The first model evaluates trips generation, focusing on trips originating from public transport stations, while the second evaluates trips attraction, analyzing trips ending at public transport stations. These models were chosen because public transportation trips often exhibit significant variability and over-dispersion, where the variance exceeds the mean, making traditional Poisson regression unsuitable for our dataset. The NB models effectively address over-dispersed count data, allowing us to better capture the variability in public transportation trips across different zones and providing a more accurate understanding of the factors influencing both trip generation and attraction scenarios.
@@
The choice of NB models aligns with previous studies that have employed similar approaches to analyze transportation demand across various modes (e.g., Schimohr and Scheiner, 2021; Hu et al., 2022; Manville et al., 2023). It is also consistent with prior research examining behavioral and spatial patterns in Santiago (Graells-Garrido et al., 2017; Beiró et al., 2018; Graells-Garrido et al., 2023b).
On the other hand, the DiD methodology estimates causal effects by comparing changes in outcomes over time between treatment and control groups before and after an intervention. In this study, the treatment group comprises Traffic Analysis Zones (TAZs) where shared e-scooter trips were observed, while the control group includes zones without e-scooter trips.
The DiD design assumes parallel trends: in the absence of shared e-scooter implementation, the treatment and control zones would exhibit similar trends in public transport demand. This assumption allows us to isolate the impact of the introduction of e-scooter from other factors that vary over time. The following equations represent the DiD structure incorporated into our NB regression models. Specifically, we employ two models: one focused on trip generation, and another on trip attraction. These models enable a comprehensive understanding of the effects of e-scooters on public transport demand.
The distinction between treatment and control zones, as detailed in the section Clustering and Definition of Treatment and Control Groups, is important to ensure the validity of the difference-in-differences (DiD) approach. Treatment zones were defined as those where shared e-scooter trips occurred, while control zones were selected based on their lack of e-scooter trips and their location outside a 1,440-meter buffer zone. This approach minimizes the likelihood of indirect influences from nearby e-scooter activity on control zones. By applying this distinction within the regional clusters defined earlier, we ensured that comparisons between treatment and control zones reflect differences attributable to e-scooter introduction, rather than variations in sociodemographic or transport characteristics.
For the trip generation model, we examine how e-scooters arriving at a zone (destinations) affect the generation of public transport trips from that same zone, evaluating whether e-scooters function as feeders that bring people to transit stations. Conversely, for the trip attraction model, we analyze how e-scooters departing from a zone (origins) influence public transport trips ending in that zone, assessing whether e-scooters serve as last-mile connections from transit stations.
$\ln\left({\textbf{E}}[y_{gen}\mid\mathbf{x}_{gen}]\right)=$ $\,\beta_{0}+\beta_{1}\text{treatment\_zones}+\beta_{2}\text{post}$ (1)
[…]
$+\mathbf{\beta}\cdot\text{Controls}+\epsilon,$
$\ln\left({\textbf{E}}[y_{att}\mid\mathbf{x}_{att}]\right)=$ $\,\beta_{0}+\beta_{1}\text{treatment\_zones}+\beta_{2}\text{post}$ (2)
[…]
$+\mathbf{\beta}\cdot\text{Controls}+\epsilon,$
where $y_{gen}$ and $y_{att}$ represent the count of public transport trips in a TAZ for trip generation and attraction scenarios, respectively. The variable treatment_zones is a dummy variable indicating whether the zone is associated with e-scooter trips: for the trip generation model, it takes the value 1 if e-scooter trips end in the zone (destinations), and for the trip attraction model, it takes the value 1 if e-scooter trips start in the zone (origins). The variable post is a dummy indicating the post-implementation period (2019), and the interaction term treatment×post captures the difference-in-differences (DiD) effect.
The $\mathbf{\beta}\cdot\text{Controls}$ term incorporates additional control variables to account for variations in trip counts due to socio-demographic, temporal, and spatial factors. These control variables include the type of day (Monday-to-Thursday, Friday, Saturday, or Sunday, with Monday-to-Thursday as the reference category) and the time period during which trips occur (morning valley, morning peak, lunch, afternoon valley, afternoon peak, night, or night valley, with morning valley as the reference category). Spatial and demographic characteristics such as the log-transformed count of the working population in the TAZ, the average educational level relative to the regional average, and the number of metro lines within a 750-meter radius are also included. By accounting for these variables, the models ensure robustness in estimating the effects of e-scooter implementation, isolating its impact from other potential confounding factors.
@@
For each scenario (generation and attraction), we estimated three separate models: (1) a model focusing on trips whose first or last stage involves a bus; (2) a model focusing on trips whose first or last stage involves the metro; and (3) a model that includes all public transport trips that involve either bus or metro in any stage of the journey. In all models, the dependent variable is the count of public transport trips in each Traffic Analysis Zone (TAZ). We applied the regression to three regions (Central, Intermediate, and Peripheral), each containing its respective control and treatment areas.
Our analysis focused on two distinct time periods in fall: for public transport, we used data on boardings and alightings from the second week of May in both 2018 and 2019; for scooters, we calculated the average number of trips between March 1 and April 15, 2019. Ideally, we would have used data for both modes from the same week—either in April or May—but this was not possible due to data availability. We therefore assumed that e-scooter usage remained relatively stable from the March–April period to mid-May.
A difference-in-differences approach was utilized to control for temporal trends and other unobservable differences between the zones that implemented scooters (treatment group) and those that did not (control group).
@@
For the difference-in-differences design, we established a minimum threshold of 5 daily e-scooter trips per zone as the treatment assignment criterion. This threshold ensures that only zones with sufficient e-scooter activity are considered ”treated” by the intervention. Based on empirical analysis of zone-day observations for both e-scooter origins (403 observations) and destinations (420 observations), the 5-trip threshold is positioned between the 25th percentile (3.0 trips/day for both) and the median (8.0 trips/day for origins, 8.75 trips/day for destinations). This results in 64.3% of origin observations and 66.9% of destination observations being assigned to treatment, with consistent assignment rates across flow types (2.6 percentage point difference). The threshold effectively excludes zones with sporadic activity that would not be meaningfully exposed to e-scooter interventions, preserving the internal validity of the difference-in-differences design.
@@ 5 Results
In this section, we present our findings on the impact of shared e-scooter introduction on public transportation usage in Santiago, Chile. Using negative binomial regression with a difference-in-differences approach, we analyze how the deployment of e-scooter services affects public transport demand across urban regions. Our analysis distinguishes between trip generation (public transport boardings - trips originating in each zone) and attraction (public transport alightings - trips ending in each zone) to capture the multidirectional effects of this new micro-mobility service.
@@
Tables in Appendix C present the results of the negative binomial models for public transport trip generation. Table 3 displays the marginal effects of the introduction of shared e-scooters on public transport trip generation across different regions. Table 4 shows the average public transport trips by region and mode of transport, providing baseline context for interpreting the percentage changes in trip generation following e-scooter introduction.
We examine how e-scooter trips arriving at each zone influence the generation of public transport trips from those same zones. This analysis reveals whether e-scooters function as feeders to public transport (bringing people to transit stations who then continue their journeys via bus or metro) or whether they operate independently. Our findings demonstrate spatially heterogeneous effects across regions and transport modes.
In the Central Region, the introduction of shared e-scooters is associated with a statistically significant reduction in total public transport boardings of 465.81 trips per hour per zone, representing a 23.87% decrease. When examining individual transport modes separately, neither bus boardings (20.93 additional trips per hour per zone, 11.01% increase) nor metro boardings (450.92 additional trips per hour per zone, 10.17% increase) show statistically significant changes.
In the Intermediate Region, the introduction of shared e-scooters shows contrasting effects across transport modes. Total public transport boardings increased significantly by 178.99 trips per hour per zone, corresponding to a 33.64% increase. Metro boardings alone also experienced a statistically significant increase of 210.53 trips per hour per zone (9.77% increase). However, bus boardings showed a minimal and non-significant decrease of 0.53 trips per hour per zone (-0.59%).
In the Peripheral Region, no statistically significant effects were observed across any transport mode. Metro boardings could not be evaluated for this region due to the limited number of metro stations.
Mode of Transport Central Intermediate Peripheral
Bus and Metro -465.81* 178.99* 52.43
(-23.87%*) (33.64%*) (18.23%)
Bus 20.93 -0.53 4.30
(11.01%) (-0.59%) (4.99%)
Metro 450.92 210.53*
(10.17%) (9.77%*)
Table 3: Difference-in-differences estimates of e-scooter impact on public transport trip generation (boardings). Marginal effects are expressed as additional trips per hour per zone, with percentage changes relative to baseline levels (Table 4) shown in parentheses. Estimates are derived from negative binomial regression models controlling for temporal, spatial, and demographic factors. Results in bold with asterisks indicate statistical significance at the 5% level ($p<0.05$).
Region Boardings (trips/hr/zone) Alightings (trips/hr/zone)
Bus Metro Bus or Bus Metro Bus or
Metro Metro
Central 190.09 4435.24 1951.40 134.21 1186.10 540.08
Intermediate 89.26 2155.23 531.86 62.78 460.68 136.52
Peripheral 86.04 287.60 49.97 71.12
Table 4: Baseline public transport demand by region and mode. Values represent average trips per hour per zone during the pre-treatment period (May 2018), providing context for interpreting percentage changes in Tables 3 and 5. Boardings capture trip origins (generation model) while alightings capture trip destinations (attraction model).
@@
Tables in Appendix D present the results of difference-in-differences models for public transport trip attraction, while Table 5 shows the marginal effects of e-scooter introduction on public transport trips ending in different urban regions. The table includes both absolute values (additional trips per hour per zone) and percentage changes in parentheses, allowing for a comprehensive assessment of the impact across all transport modes and regions.
Mode of Transport Central Intermediate Peripheral
Bus and Metro -15.58 5.57* 3.07
(-2.88%) (4.08%*) (4.32%)
Bus 13.88 1.31 4.25
(10.34%) (2.09%) (8.51%)
Metro 9.19 -35.52
(0.77%) (-7.71%)
Table 5: Difference-in-differences estimates of e-scooter impact on public transport trip attraction (alightings). Values represent marginal effects in additional trips per hour per zone, with percentage changes relative to baseline levels (Table 4) in parentheses. Results from negative binomial regression models controlling for temporal, spatial, and demographic factors. Values in bold and with asterisks are statistically significant at $p<0.05$.
Regarding the impact of e-scooter introduction on public transport alightings, the analysis reveals heterogeneous effects across regions and transport modes. For total public transport alightings (bus and metro), the Intermediate Region showed the only statistically significant effect with an increase of 5.57 trips per hour per zone (4.08%), while the Central Region showed a non-significant decrease (-15.58 trips per hour per zone, -2.88%) and the Peripheral Region showed a non-significant increase (3.07 trips per hour per zone, 4.32%). For individual transport modes, all effects were non-significant. Bus alightings demonstrated increases across all regions (Central: 13.88 trips per hour per zone, 10.34%; Intermediate: 1.31 trips per hour per zone, 2.09%; Peripheral: 4.25 trips per hour per zone, 8.51%), while metro alightings exhibited mixed patterns with an increase in the Central Region (9.19 trips per hour per zone, 0.77%) and a decrease in the Intermediate Region (-35.52 trips per hour per zone, -7.71%). Metro alightings could not be evaluated in the Peripheral Region due to limited metro infrastructure coverage.
5.1 Robustness and Model Validation
We examine the robustness of our difference-in-differences estimates by comparing our baseline specification with an extended model that includes demographic controls. This analysis serves two purposes: first, it assesses the stability of treatment effects when controlling for observable characteristics that may correlate with both treatment assignment and travel outcomes; second, it provides insights into the mechanisms through which the intervention affects travel behavior across different urban areas.
Our baseline specification (Model 1) includes the core DiD variables (treatment zones, post-period indicator, and their interaction), metro accessibility, and type of day and time period fixed effects. The extended specification (Model 2) augments this baseline with additional demographic controls: log working population and relative educational level. We examine these specifications across three distinct geographic areas that capture different urban contexts: Central Region, Intermediate Region, and Peripheral Region.
Tables 6 and 7 present the results for trip generation and attraction models, respectively, across the three geographic regions: Central, Intermediate, and Peripheral. Both specifications include metro accessibility as a fundamental control for public transit infrastructure. Model 2 extends the baseline by adding demographic controls (log working population and relative educational level) to assess the robustness of treatment effects to observable characteristics that may correlate with both treatment assignment and travel outcomes.
For trip generation, treatment effects remain statistically significant and economically meaningful in both the baseline model (without sociodemographic controls) and the extended model (with controls) for central and intermediate areas. In central areas, the negative treatment effect decreases modestly from -0.282 to -0.249 when controls are introduced, but retains significance at the 5% level. This suggests successful modal substitution or trip consolidation in areas with high existing accessibility. In intermediate areas, the positive and significant effect declines from 0.413 to 0.358, indicating induced demand or improved connectivity. In contrast, the effect observed in the peripheral region is sensitive to the inclusion of sociodemographic controls: the coefficient decreases from 0.224 to 0.185 and loses statistical significance, suggesting that socioeconomic characteristics play a key mediating role in these more remote locations.
Variable Central Intermediate Peripheral
Model 1: Baseline DiD + Metro
Treatment $\times$ Post -0.2818* 0.4134* 0.2241*
(0.055) (0.094) (0.088)
Metro 0.657* 0.075 -0.027
(0.142) (0.055) (0.079)
R-squared 0.024 0.031 0.031
Model 2: DiD + Metro + Demographic Controls
Treatment $\times$ Post -0.2491* 0.3583* 0.1854
(0.052) (0.089) (0.091)
Metro 0.229* 0.178* 0.055
(0.041) (0.033) (0.052)
Log working population 0.352* 0.356* 0.198*
(0.065) (0.078) (0.089)
Relative education level 0.688* 0.374* 0.249*
(0.112) (0.095) (0.108)
R-squared 0.036 0.037 0.034
Table 6: Robustness analysis for public transport trip generation models across geographic regions. Values represent regression coefficients from negative binomial models, with standard errors clustered at the zone level in parentheses. Model 1 includes core difference-in-differences variables, metro accessibility, and temporal fixed effects. Model 2 adds demographic controls (log working population and relative educational level). Values in bold with asterisks are statistically significant at $p<0.05$.
The robustness analysis of total public transportation usage (metro and bus boardings and alightings) reveals several important findings. First, the inclusion of demographic controls consistently improves model fit across regions and specifications, as evidenced by increases in R-squared values. This improvement indicates that demographic characteristics capture meaningful variation in travel behavior that would otherwise remain in the error term.
Variable Central Intermediate Peripheral
Model 1: Baseline DiD + Metro
Treatment $\times$ Post 0.0054 0.0276 0.0486
(0.043) (0.039) (0.067)
Metro access 0.530* 0.049 0.115
(0.089) (0.041) (0.067)
R-squared 0.029 0.025 0.024
Model 2: DiD + Metro + Demographic Controls
Treatment $\times$ Post -0.0306 0.0418* 0.0436
(0.041) (0.037) (0.065)
Metro access 0.132* 0.103* 0.179*
(0.035) (0.028) (0.045)
Log working population 0.308* 0.281* 0.075*
(0.058) (0.062) (0.071)
Relative education level 0.630* 0.341* 0.168*
(0.098) (0.085) (0.095)
R-squared 0.039 0.030 0.025
Table 7: Robustness analysis for public transport trip attraction models across geographic regions. Values represent regression coefficients from negative binomial models, with standard errors clustered at the zone level in parentheses. Model 1 includes core difference-in-differences variables, metro accessibility, and temporal fixed effects. Model 2 adds demographic controls (log working population and relative educational level). Values in bold with asterisks are statistically significant at $p<0.05$.
Second, the comparison between trip generation and attraction models reveals a coherent asymmetry that reflects e-scooter adoption patterns, particularly in the intermediate region. Treatment effects are substantially larger and more robust for trip generation than for trip attraction, consistent with e-scooters functioning primarily as first-mile connectivity solutions in intermediate areas, where users predominantly employ these devices to access public transportation nodes rather than for last-mile distribution from transit stops to final destinations. This contrasts with the central region, where e-scooters appear to substitute for public transportation in trip generation, while showing no significant impact on trip attraction patterns.
Third, the persistence of significant treatment effects in the presence of demographic controls demonstrates the robustness of our findings. The fact that estimated effects remain statistically significant and of substantial magnitude after controlling for observable characteristics that could correlate with treatment assignment strengthens the causal interpretation of our results. The modest changes in coefficient magnitudes suggest that our treatment effects are not merely reflecting unobserved heterogeneity in demographic composition across treatment and control areas.
The consistency of our results across specifications and geographic areas provides strong evidence for the robustness of our causal estimates. Treatment effects maintain their magnitude and significance when including demographic controls, while the distinct spatial patterns observed across central and intermediate areas validate that we are capturing genuine behavioral responses to e-scooter availability rather than confounding factors.